How do I answer this?: .75=x-275/80
Answer by Theo(13342)

your equation is:
.75=x-275/80
add 275/80 to both sides of this equation to get:
.75 + 275/80 = x
commute this to get:
x = .75 + 275/80
you can use your calculator to solve for x to get:
x = .75 + 3.43754 which becomes:
x = 4.1875
alternatively, you can keep it in fraction form as follows:
.75 is the same as 75/100
your equation becomes:
75/100 + 275/80 = x
commute this to get:
x = 75/100 + 275/80
put these under a common denominator of (100*80) to get:
x = ((75*80) + (275*100)) / (100*80)
simplify to get:
x = (6000 + 27500)/8000
simplify further to get:
x = 33500/8000
divide numerator and denominator by 100 to get:
x = 335/80
divide numberator and denominator by 5 to get:
x = 67/16
i don't think it can be simplified further than that.
that would be your answer in fraction form.
divide 67/16 and you get:
x = 4.1875
since that's the same answer we got up top, we did the fraction form correctly.
how you provide the answer is determined by what your instructor expects.
either form is correct.
x = 4.1875 or x = 67/16 are your answers to this problem.